You are building a new apartment building and need to decide between installing plastic waste pipe at $100,000 or cast iron pipe at $120,000. The plastic pipe will need partial replacement in 15 years, which will cost $70,000. The cast iron pipe will cost nothing extra for at least 30 years. Using an APR of 10%, determine which is more economical.
Answer #1

Plastic Waste Pipe:

Initial cost = $100000

Replacement at end of 15 years = $70000

PV of replacement cost = 70000/(1+10%)^15 = 70000/1.1^15 = 70000/4.1772 = $16757.44

Hence Total PV of cost = 100000+16757.44 = $116757.44

Cast Iron Pipe Initial cost = $120000.

PV of cost =$120000 since no replacement required

Since PV of Plastic Waste pipe is lower it is economical